TotsuGami00:

A manufacturing unit currently operates at 80 percent of its capacity. The profit function for the unit at the optimum output, x, is given by p(x) = -0.1x2 + 80x − 60.

AngeI:

So what are you doing with this equation?

TotsuGami00:

If the function f(x) models the current capacity of the unit, the composite function giving the unit's current profit function is . If the optimum output is 500 units, the current profit is $.

AngeI:

I'm sorry if I'm wrong but wouldn't you just fill in 500 for x?

TotsuGami00:

it's some p fx crap or f px for the first

AngeI:

Oh okiii Also i forgot to ask, is there answer choices?

TotsuGami00:

f(p(x))=-0.064x²+6.4x-60 p(f(x))=-0.064x²+64x-60 f(p(x))=-0.064x²+64x-60 p(f(x))=-0.64x²+6.4x-60 p(f(x))=-0.64x²+64x-60

TotsuGami00:

For the 2nd box, 15,400 15,940 16,060 16,600

AngeI:

I believe it's b? For both e.e I'm sorry I can't really explain v.v Also it looks like this question was asked before but its a mess e.e https://questioncove.com/updates/54518467e4b0be646b99cf7e

TotsuGami00:

k. thnks